1a) At the school play on Fri. night 120 tickets were sold. Student tickets cost $5 each, and Adult tickets colst $7 each. the total sale for the evening was $670. How many of each type of tickets were sold?
1b) Write a sytem of equations that represents this problem, then solve and asnwer the question.
Answer for 1a- amount of Student tickets sold is 85
amount of Adult tickets sold is 35
Could you please help me with 1b I'am not exactly sure how to do it :(

Answer by scott8148(6628)

let x=students and y=adults ___ x+y=120 ___ x=120-y

5x+7y=670 ___ 5(120-y)+7y=670 ___ 600-5y+7y=670 ___ 2y=70